It took an airplane 2hours to fly 600mi against a headwind. The return trip with the wind took 1 2/3hours. Find the speed of the plane with no wind and the windspeed.
The speed of the plane is mph.
Windspeed mph.


Answer by TimothyLamb(4379)

s = d / t
---
(a - w) = 600 / 2
a - w = 300
a = 300 + w
---
(a + w) = 600 / (1+(2/3))
(a + w) = 600 / ((3/3)+(2/3))
(a + w) = 600 / (5/3)
a + w = 360
---
a + w = 360
300 + w + w = 360
2w = 60
---
w = 30 mph
a = 330 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php